OpenStudy (idealist10):

Find all solutions of (sin x)(sin y)+(cos y)y'=0.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

Looks separable, doesn't it?

OpenStudy (anonymous):

\[\begin{align*} \sin x\sin y+\cos y~y'&=0\\ \cos y~dy&=-\sin x\sin y~dx\\ \cot y~dy&=-\sin x~dx \end{align*}\]

OpenStudy (idealist10):

Wait a minute. Let me work it out.

OpenStudy (anonymous):

If you're not familiar with the integral result of \(\displaystyle\int\cot x~dx\), consider it as \(\displaystyle\int\frac{\cos x}{\sin x}~dx\), then use a substitution.

OpenStudy (idealist10):

I'm familiar, it's ln abs(sin y)=cos x+C, right? But how do I solve for y?

OpenStudy (anonymous):

Yes that's correct. Take the exponential of both sides: \[\large e^{\ln|\sin y|}=e^{\cos x+C}~~\iff~~\sin y=e^Ce^{\cos x}~~\iff~~\sin y=Ce^{\cos x}\] Then take the inverse sine.

OpenStudy (idealist10):

so the answer is y=sin^-1 (ce^cos x), right?

OpenStudy (anonymous):

Yes I think so. Usually, an implicit solution is acceptable. I would have stopped at \(\sin y=\cdots\), but stick to your instructions.

OpenStudy (idealist10):

Okay, thank you!

OpenStudy (anonymous):

You're welcome! Checking with Wolfram: http://www.wolframalpha.com/input/?i=Sin%5Bx%5D*Sin%5By%5D%2BCos%5By%5D*y%27%3D0 They skip the step where you have \(e^C=C\), but it's still right.